-3x + 4y= 8
x +2y =10

Answer by oberobic(2304)

-3x + 4y = 8
x + 2y = 10
.
Multiply the second equation by 2 and then subtract the second equation to eliminate the 'y' term.
.
-3x +4y = 8
2x + 4y = 20
------------- subtract
-5x = -12
x = -12/-5 = 12/5
.
Now substitute to find 'y' using the first equation.
.
-3*(12/5) + 4*y = 8
4y = 8 +36/5 = 40/5 + 36/5 = 76/5
y = 19/5
.
Check these values using the second equation to be sure this is the answer.
(12/5) + 2*(19/5) = 10
12/5 + 38/5 = 50/5 = 10
Correct.
.
Answer: x = 12/5, y = 19/5, which can be shown as the ordered point (12/5, 19/5).
.
Done.